OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y = 4/9 X – 2 that passes through the point (4,3)?

OpenStudy (anonymous):

in this ques equation y = 4/9x-2 is given which is of the form of y = m1x+c where m1 is a slope on comparing with this we can find out the slope now in ques, we need to find the equation of line which is perpendicular to the given line from this it is concluded that slope of both the line would be perpendicular to each other i.e. m1*m2 = -1 from this slope of required line can be found out after this u need to find the equation which can be found out by using the formula y-Y1 = m2(x-X1) just put the values in above formula u will get the required equation of line

OpenStudy (anonymous):

still a bit confused.

OpenStudy (anonymous):

in which step u r confused

OpenStudy (anonymous):

The last part

OpenStudy (anonymous):

y-Y1 = m2(x-X1)

OpenStudy (anonymous):

in the last part we r going to find the equation of line passing through the point (4,3) from above we can conclude x1 = 4 and y1=3 u know the m2 which has been found in the previous step stated above u need to put values in the formula y-Y1= m2(x-X1)

OpenStudy (anonymous):

@asap1 is that right???????

OpenStudy (anonymous):

can i just get the answer?

OpenStudy (anonymous):

i m getting equation of line 4y = -9x+48 y = -(9/4)x+12

OpenStudy (anonymous):

@asap1 is this answer right????????

OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y = 4/9x – 2 that passes through the point (4,3). y = mx + b m is slope b is y-intercept ----------------------- Slope of perpendicular lines are opposite reciprocal. The slope of the given line is 4/9. The opposite reciprocal of 4/9 is -9/4. So the slope of the new line is -9/4. So far you have: y = -9/4x ------------------------ Now you need to find the value of the y-intercept. To do this, put in the values of the given point for x and y. (x, y) (4, 3) y = mx + b 4 = -9/4(3) + b Solve for b and plug it into the equation: y = -9/4x + b